OpenStudy (anonymous):

Let (theta) = 2pi/3. Find the exact values of csc(theta), sin(theta), and cot(theta).

OpenStudy (dls):

\[\LARGE \sin \frac{2\pi}{3}=\sin(120^{0})=\sin(90^0+30^0)=?\]

OpenStudy (anonymous):

idk

OpenStudy (dls):

sin(90+theta)=cos theta cos30=?

OpenStudy (anonymous):

still don't know

OpenStudy (anonymous):

Look at this: http://www.mathsisfun.com/geometry/images/circle-unit-304560.gif Pay close attention to the top right "legend" It shows that the x-coordinate is cos and the y-coordinate is sin. Since Θ = (2π/3) or 120° So \[sin(Θ) = (√3)/2\] And \[cos(Θ) = (-1/2)\] Check that with the link i sent you. This question asks for csc(Θ), sin(Θ), cot(Θ). I'll start with showing you how that works once you give me the all-clear that you see where i'm going so far.

OpenStudy (anonymous):

all clear!

OpenStudy (anonymous):

I'll start with csc(Θ) csc(Θ) = 1/cos(Θ) From the previous work I mentioned that the cos(120) = -(1.2) So: \[\csc (120) = \frac{ 1 }{ \cos (120)} \] \[\csc (120) = \frac{ 1 }{ -\frac{1}{2}} \]